# Towards Understanding Domain Adapted Sentence Embeddings for Document Retrieval

### v2

@@ 4.2 Isotropy Score Analysis
sciq
bge_large 92.7 9.29, 82.99 93.2 9.58, 97.89 92.7 9.73, 82.77 89.4 30.41, 91.88
bge_large_pretrained 92.1 5.04, 85.08 92.6 5.29, 97.05 91.9 5.31, 85.13 86.4 31.95, 91.82
bge_large_finetuned 94.5 5.77, 73.93 94.8 6.21, 97.26 94.8 6.18, 73.94 89.8 24.63, 91.84
bge_large_pretrained_finetuned 95.5 6.83, 76.76 94.7 7.22, 97.51 95.5 7.21, 76.83 89.4 28.32, 91.82
llm_embedder 91.7 10.62, 67.87 93.0 10.96, 96.77 91.1 11.8, 66.38 86.9 29.65, 91.24
llm_embedder_pretrained 91.0 6.43, 73.75 91.3 6.78, 96.1 91.0 6.8, 72 83.7 25.18, 91.13
llm_embedder_finetuned 94.4 8.49, 65.9 93.9 9.06, 96.51 94.1 9.05, 65.25 86.8 28.81, 91.15
llm_embedder_pretrained_finetuned 94.0 10.08, 67.46 93.5 10.48, 96.92 93.5 10.57, 66.48 86.5 27.02, 91.13
PubMedQA
bge_large 94.0 11.09, 79.17 95.8 11.59, 97.83 93.7 11.48, 79 91.5 26.6, 92.78
bge_large_pretrained 84.3 4.96, 76.3 81.1 5.39, 95.8 81.1 84.07, 76.44 80.2 27.01, 92.76
bge_large_finetuned 98.5 8.25, 72.33 97.3 8.76, 97.18 98.2 8.76, 71.56 93.4 27.3, 92.77
bge_large_pretrained_finetuned 97.8 6.42, 72.42 96.8 6.92, 96.69 98.2 6.88, 72.52 94.0 27, 92.74
llm_embedder 95.4 14.93, 63.95 95.3 15.67, 96.75 95.3 15.6, 62.68 90.6 33.86, 90.33
llm_embedder_pretrained 91.9 6.41, 68.04 90.7 7.09, 94.19 91.4 7.08, 67.35 85.7 28.7, 90.19
llm_embedder_finetuned 97.5 12.81, 63.51 97.5 13.53, 96.42 97.9 13.53, 62.33 92.6 33.97, 90.39
llm_embedder_pretrained_finetuned 97.5 8.96, 63.96 97.1 9.77, 95.59 97.0 9.67, 62.5 91.8 32.72, 90.35
Table 5: Accuracy, $I_{A}$ and $I_{B}$ for embeddings under different transformations.
Combining these observations, we conclude that fine tuning improves the isotropy but isotropy cannot be attributed to retrieval accuracies. Our studies indicate that this may be the right resolution between the contradictions among studies by [12] and [13] which we have discussed in Section 1.
4.3 Embedding Analysis post Domain Adaptation
As described in Section 2.6, Figure 2 shows the distribution of distances $\Delta_{\mathcal{M}}$ and $\Delta_{\mathcal{M^{\prime}}}$ for both bge_large and llm_embedder, respectively. The plots show that post domain-adaptation, the embeddings move away from the domain-agnostic embeddings. For llm_embedder, the farthest points between $\mathcal{D}$ and $\mathcal{D^{\prime}}$ on public models is closer than the closest ones post domain adaptation. For bge_large, it is clear that distance between embeddings increase post domain adaptation, but this is less pronounced on Sci-Q and PubMedQA datasets. This is one of the effects of domain-adaption and needs further study.
5 Recommendations and Conclusions
5.1 Recommendations
@@
  • If possible, use domain adaptation - preferably pretraining followed by fine-tuning (PTFT)
  • Identify thresholds for the similarity scores - this will lead to bootstrapped accuracy which is statistically same as the full dataset bootstrapped accuracy, while suppressing less relevant documents to end-users / downstream tasks
@@
5.2 Conclusion and Future Work
@@
We have reported mean bootstrapped retrieval accuracies along with confidence intervals for various SOTA embedding models with and without domain-adaptations. We observe that fine-tuning (with or without pre-training) improves both mean and CI of retrieval accuracies. However, pre-training followed by fine-tuning improves CI further. We propose a bootstrapped approach for choosing thresholds and observe that we can significantly reduce the number of retrieved sentences without any statistical deviation in retrieval performance. Our proposed cumulative distribution metrics, COE and ROE, to measure overlap between distributions of cosine similarities show strong correlations with retrieval performance and similarity thresholds respectively. We measure isotropy of embeddings using two independent SOTA isotropy metrics. We perform extensive evaluations on embeddings with and without isotropic transformations. We conclude that isotropy can be considered to be neither necessary nor sufficient from a retrieval accuracy perspective. Finally, we observe that with domain adaption, domain specific embedding show improved isotropy scores and move away from general domain embeddings. Our study establishes systematic methods of analysing embeddings in specialised domains. Our results hold across three different domain, which makes us believe that they hold for other specialised domains too. The current work considers QA task only. Future work may involve other tasks like summarization, or multimodal settings.
(a) bge-large-en, TeleQuAD
(b) bge-large-en, PubMedQA
(c) bge-large-en, sci-q
(d) llm-embedder, TeleQuAD
(e) llm-embedder, PubMedQA
(f) llm-embedder, sci-q
Figure 2: Distribution of distances for embedding models. Blue histogram represents $\Delta_{\mathcal{M}}$ and orange represents $\Delta_{\mathcal{M}^{\prime}}$ using bge_large and llm_embedder models as $\mathcal{M}$. Refer Sec. 2.6 for definitions
Appendix A Appendix
A.1 Detailed metrics for various datasets
Embedding Model Bootstrapped Acc (CI) COE (CI) ROE (CI) $\tau,\psi$ Acc @ $\tau$
Dataset: TeleQuAD, k=5
gte 72.55 (64.48, 81.0) 87.66 (83.08, 92.35) 0.31 (0.20, 1.02) 0.28 (20) 69.75
bge_large 66.87 (58.48, 75.52) 87.98 (82.06, 94.70) 4.81 (1.22, 18.04) 0.5 (35) 67.18
bge_large_pretrained 62.64 (53.0, 70.0) 85.94 (79.92, 93.68) 2.18 (0.61, 6.52) 0.58 (25) 61.36
bge_large_finetuned 81.61 (74.48, 88.52) 91.98 (90.42, 94.09) 0.22 (0.00, 0.82) 0.43 (25) 79.46
bge_large_pretrained_finetuned 81.67 (74.48, 87.52) 91.06 (89.81, 93.07) 0.23 (0.20, 0.31) 0.4 (35) 77.73
llm_embedder 70.06 (63.0, 77.52) 87.26 (80.63, 96.13) 5.77 (0.92, 26.71) 0.78 (30) 69.9
llm_embedder_pretrained 57.12 (47.48, 67.05) 84.88 (79.71, 91.34) 6.32 (1.94, 19.27) 0.75 (30) 52.53
llm_embedder_finetuned 81.58 (75.0, 88.52) 90.73 (88.38, 95.11) 0.10 (0.00, 0.82) 0.56 (40) 80.69
llm_embedder_pretrained_finetuned 80.37 (74.0, 86.52) 90.74 (88.99, 94.70) 0.21 (0.10, 0.41) 0.53 (25) 77.97
ada_002 75.48 (68.47, 83.0) 90.19 (85.02, 95.92) 3.21 (1.33, 9.28) 0.75 (25) 75.3
text-embedding-3-small 69.91 (62.0, 79.0) 87.59 (83.38, 95.01) 1.90 (0.61, 8.05) 0.31 (20) 68.43
text-embedding-3-large 75.96 (69.0, 82.0) 90.32 (86.03, 96.74) 5.99 (2.14, 23.75) 0.26 (20) 73.72
mpnet 61.49 (54.0, 71.0) 81.74 (75.54, 91.34) 2.91 (0.82, 11.21) 0.29 (45) 59.78
minilm 67.26 (60.0, 75.0) 83.25 (78.90, 90.72) 0.77 (0.20, 3.47) 0.27 (25) 64.98
Dataset: sci-qa, k=5
bge_large 92.7 (87, 97) 96.98 (95.85, 99.39) 4.14 (2.2, 13.9) 0.508 (85) 92.8
bge_large_pretrained 92.08 (87, 96) 95.48 (94.02, 97.8) 1.99 (1.34, 4.39) 0.5 (15) 92.1
bge_large_finetuned 94.45 (90, 98) 97.79 (96.95, 99.27) 3.72 (2.07, 8.54) 0.883 (85) 94.8
bge_large_pretrained_finetuned 95.5 (91, 98) 98.15 (97.2, 99.27) 3.25 (1.71, 8.05) 0.771 (85) 95.4
llm_embedder 91.69 (87, 96) 97.2 (95.49, 99.27) 6.85 (2.93, 22.68) 0.809 (85) 91.8
llm_embedder_pretrained 91.03 (86, 96) 94.42 (92.56, 98.17) 2.21 (1.22, 5.61) 0.682 (85) 91.2
llm_embedder_finetuned 94.37 (90, 98) 98.18 (97.32, 99.51) 7.4 (3.17, 24.39) 0.883 (85) 94.7
llm_embedder_pretrained_finetuned 94 (90, 98) 98.24 (97, 99.51) 6.28 (2.32, 23.05) 0.824 (85) 93.9
Dataset: PubMedQA, k=5
bge_large 94 (89, 98) 96.2 (95.1, 98.4) 3.21 (1.3, 13.1) 0.426 (5) 93.8
bge_large_pretrained 84.3 (78, 90) 93.8 (90.3, 99) 7.52 (3.2, 23.8) 0.492 (5) 83.4
bge_large_finetuned 98.51 (95, 100) 99.5 (99.3, 99.7) 4.81 (2.8, 10.1) 0.791 (85) 98.5
bge_large_pretrained_finetuned 97.83 (96, 100) 98.95 (98.4, 99.6) 6.71 (3.2, 21) 0.755 (45) 97.7
llm_embedder 95.42 (91, 99) 97.7 (97, 99.5) 4.49 (1, 24.1) 0.732 (5) 95.8
llm_embedder_pretrained 91.87 (86, 96) 96.94 (95.3, 99.2) 4.1 (1.6, 18.7) 0 (100) 92.0
llm_embedder_finetuned 97.53 (94, 100) 99.28 (98.8, 99.8) 5.55 (2.5, 14.88) 0.797 (30) 97.5
llm_embedder_pretrained_finetuned 97.5 (95, 100) 99 (98.6, 99.7) 6.15 (3, 22.8) 0.783 (30) 97.2
Table 6: Different metrics for various datasets ($\mathcal{D}$), $K$=5
Embedding Model Baseline Standardized Whitened PCA
Acc $I_{A},I_{B}$ Acc $I_{A},I_{B}$ Acc $I_{A},I_{B}$ Acc $I_{A},I_{B}$
TelecomQuAD
gte 72.55 11.33, 42.4 72.22 11.64, 96.75 71.98 11.64,43.14 70.85 21.34, 97.49
ada_002 75.48 6.64, 25.52 68.83 7.08, 97.09 75.46 7.06.26.18 69.31 15.67, 93.8
text-embedding-3-small 69.91 6.33, 45.86 67.18 6.79, 93.14 69.52 6.72, 45.88 66.26 14.46, 89.17
text-embedding-3-large 75.96 3.64, 63.78 71.47 4.11, 94.38 74.84 4.08, 64.33 70.94 11.83, 89.12
mpnet 61.49 9.51, 36.41 57.62 10.25, 96.66 59.99 10.22, 27.68 56.39 16.35, 93.72
minilm 67.26 19.43, 26.59 63.99 20.54, 95.56 66.19 20.37,26.36 62.22 25.67, 95.32
sciq
bge_large 92.7 9.29, 82.99 93.2 9.58, 97.89 92.7 9.73, 82.77 89.4 30.41, 91.88
bge_large_pretrained 92.1 5.04, 85.08 92.6 5.29, 97.05 91.9 5.31, 85.13 86.4 31.95, 91.82
bge_large_finetuned 94.5 5.77, 73.93 94.8 6.21, 97.26 94.8 6.18, 73.94 89.8 24.63, 91.84
bge_large_pretrained_finetuned 95.5 6.83, 76.76 94.7 7.22, 97.51 95.5 7.21, 76.83 89.4 28.32, 91.82
llm_embedder 91.7 10.62, 67.87 93.0 10.96, 96.77 91.1 11.8, 66.38 86.9 29.65, 91.24
llm_embedder_pretrained 91.0 6.43, 73.75 91.3 6.78, 96.1 91.0 6.8, 72 83.7 25.18, 91.13
llm_embedder_finetuned 94.4 8.49, 65.9 93.9 9.06, 96.51 94.1 9.05, 65.25 86.8 28.81, 91.15
llm_embedder_pretrained_finetuned 94.0 10.08, 67.46 93.5 10.48, 96.92 93.5 10.57, 66.48 86.5 27.02, 91.13
PubMedQA
bge_large 94.0 11.09, 79.17 95.8 11.59, 97.83 93.7 11.48, 79 91.5 26.6, 92.78
bge_large_pretrained 84.3 4.96, 76.3 81.1 5.39, 95.8 81.1 84.07, 76.44 80.2 27.01, 92.76
bge_large_finetuned 98.5 8.25, 72.33 97.3 8.76, 97.18 98.2 8.76, 71.56 93.4 27.3, 92.77
bge_large_pretrained_finetuned 97.8 6.42, 72.42 96.8 6.92, 96.69 98.2 6.88, 72.52 94.0 27, 92.74
llm_embedder 95.4 14.93, 63.95 95.3 15.67, 96.75 95.3 15.6, 62.68 90.6 33.86, 90.33
llm_embedder_pretrained 91.9 6.41, 68.04 90.7 7.09, 94.19 91.4 7.08, 67.35 85.7 28.7, 90.19
llm_embedder_finetuned 97.5 12.81, 63.51 97.5 13.53, 96.42 97.9 13.53, 62.33 92.6 33.97, 90.39
llm_embedder_pretrained_finetuned 97.5 8.96, 63.96 97.1 9.77, 95.59 97.0 9.67, 62.5 91.8 32.72, 90.35
Table 7: Accuracy, $I_{A}$ and $I_{B}$ for embeddings under transformations, detailed results.
A.2 Sample questions from TelecomQuad
Question Answer
What does AoA stand for? AoA stands for Angle of Arrival.
Who specifies the chargeable events? Chargeable events are specified by middle tier TS.
What architecture is defined in 3GPP TS 32.240 V15.5.0? Single IMSI Architecture is defined in 3GPP TS 32.240 V15.5.0.
Table 8: Sample questions and their correct answers - based on 3GPP Release 17.
A.3 Isotropy Scores
We provide brief details of isotropy measures used in this work. According to [25], isotropy $I_{a}(v)$ can be defined by (3).
$I_{a}(v)=\frac{\min_{|c|=1}Z(c)}{\max_{|c|=1}Z(c)},$ (3)
where $Z(c)$ is a partition function as given by (4)
$Z(c)=\sum_{\forall v}exp(c^{T}v),$ (4)
such that $Z(c)$ is constant with any unit function.
If $V$ is the matrix stacked by all embedding vectors, $1_{|V|}$ be the vectors with all entries equal to one, then the second order approximation of isotropy is given by (5).
$I_{A}(\{v\})=\frac{|V|-||I_{|V|}^{T}V||+0.5\sigma_{min}^{2}}{|V|+||I_{|V|}^{T}V||+0.5\sigma_{max}^{2}},$ (5)
where $\sigma_{min}$ and $\sigma_{max}$ correspond to smallest and largest singular value of $V$, respectively.
Further, [14] define isotropy as given by (6).
$I_{B}(v)=\frac{(n-\delta(v)^{2}(n-\sqrt{n}))^{2}-n}{n(n-1)},$ (6)
where for any embedding $v\in\mathbb{R}^{n}$, $\delta(v)=\frac{||\Sigma_{D}-1||}{\sqrt{2(n-\sqrt{n})}}$, $1=(1,1,...,1)^{T}\in\mathbb{R}^{n}$, and $\Sigma_{D}\in\mathbb{R}^{n}$ is the normalized diagonal of covariance matrix of $PCA(V)$, which is the set of embeddings transformed by the first $n$ principal components.
A.4 Distribution plots
Figure 3: Density plots for health dataset. Red, green and blue indicate distribution of $S_{rand}$, $S_{corr}$ and $S_{topK}$ respectively. Refer Sec. 2.3 for definitions
Figure 4: Density plots for science dataset. Red, green and blue indicate distribution of $S_{rand}$, $S_{corr}$ and $S_{topK}$ respectively
Figure 5: Density plots for SQuAD dataset $\mathcal{D^{\prime}}$ the PTFT model is domain adapted using telecom data. Red, green and blue indicate distribution of $S_{rand}$, $S_{corr}$ and $S_{topK}$ respectively
A.5 NDCG Scores
Table 9 shows Normalized Discounted Cummulative Gain (NDCG) [31] scores for various datasets for base, PT, FT and PT-FT variants of embedding models. The columns indicate the NDCG score, lower, upper bound and width of confidence interval, as well as accuracy and NDCG on the full dataset.
We observe that the bootstrapped accuracy numbers are comparable to those on the full dataset, this validates our approach. Further, if we consider the NDCG scores, they improve with domain adaptation. NDCG scores are relatively lower because we have only one correct answer per question. However, the improvement of NDCG scores show that not only do we get tighter confidence intervals on domain adaptation, our correct answers are retrieved with better ranks.
Model Variant TeleQuAD
NDCG NDCG-LOW NDCG-HIGH CI WIDTH Full Data Acc Full Data NDCG
bge-large Base 29.6 29.4 30.0 0.6 66.0 29.9
PT 27.2 27.0 27.4 0.4 63.1 27.5
FT 34.2 33.2 34.3 1.2 82.0 34.2
PT-FT 34.9 34.8 35.4 0.5 81.5 34.9
LLM-embedder Base 29.2 27.9 29.5 1.6 69.2 29.3
PT 25.2 24.6 25.3 0.8 57.0 25.2
FT 34.3 33.8 34.4 0.6 81.8 34.4
PT-FT 33.7 33.3 33.8 0.5 80.8 33.8
PubMedQA
NDCG NDCG-LOW NDCG-HIGH CI WIDTH Full Data Acc Full Data NDCG
bge-large Base 37.2 37.2 37.3 0.1 94.0 37.2
PT 35.1 34.9 35.2 0.4 84.3 35.0
FT 38.5 38.3 38.5 0.3 97.8 38.5
PT-FT 38.9 38.9 38.9 0.1 98.5 38.9
LLM-embedder Base 37.5 37.4 37.7 0.3 95.4 37.5
PT 37.0 36.8 37.2 0.4 91.8 37.0
FT 38.1 38.1 38.2 0.1 97.6 38.1
PT-FT 38.1 37.9 38.2 0.3 97.3 38.1
Sci-QA
NDCG NDCG-LOW NDCG-HIGH CI WIDTH Full Data Acc Full Data NDCG
bge-large Base 36.8 36.8 36.9 0.1 92.8 36.9
PT 36.7 36.6 36.8 0.2 92.1 36.7
FT 37.4 37.2 38.5 1.4 94.5 37.5
PT-FT 37.8 37.7 37.9 0.2 95.4 37.9
LLM-embedder Base 36.9 36.8 37.2 0.4 91.8 36.8
PT 36.3 36.1 36.4 0.3 91.0 36.3
FT 37.7 37.6 37.9 0.3 94.0 37.6
PT-FT 37.7 37.5 37.7 0.1 94.0 37.7
Table 9: NDCG Scores for various datasets.

### v3

@@ IV-B Isotropy Score Analysis
TABLE III: Accuracy, $I_{A}$ and $I_{B}$ for embeddings under different transformations.
Correlation of $I_{A}$ and $I_{B}$ with accuracies across base, fine-tuned models with and without post-processing using transformations described in Section II-E is presented in Table I. We see that, accuracy and both the isotropy scores are not correlated across datasets. Contrary to the conflicting claims in [12] and [13], our experiments establish that accuracy and isotropy scores are not correlated.
Combining these observations, we conclude that fine tuning improves the isotropy but isotropy cannot be attributed to retrieval accuracies. Our studies indicate that this may be the right resolution between the contradictions among studies by [12] and [13] which we have discussed in Section I.
V Recommendations and Conclusions
V-A Recommendations
@@
  • If possible, use domain adaptation - preferably pretraining followed by fine-tuning (PT-FT).
  • Identify thresholds for the similarity scores - this will lead to bootstrapped accuracy which is statistically same as the full dataset bootstrapped accuracy, while suppressing less relevant documents to end-users / downstream tasks.
@@
V-B Conclusions and Future Work
@@
We have reported mean bootstrapped retrieval accuracies along with confidence intervals for various SOTA embedding models with and without domain-adaptation. We observe that fine-tuning (with or without pre-training) improves both mean and CI of retrieval accuracies. However, pre-training followed by fine-tuning improves CI further. We proposed a bootstrapped approach for choosing thresholds and observe that we can significantly reduce the number of retrieved sentences without any statistical deviation in retrieval performance. Our proposed cumulative distribution metrics, COE and ROE, to measure overlap between distributions of cosine similarities show strong correlations with retrieval performance and similarity thresholds respectively. We measure isotropy of embeddings using two independent SOTA isotropy metrics. We perform extensive evaluations on embeddings with and without isotropic transformations. We conclude that isotropy can be considered to be neither necessary nor sufficient from a retrieval accuracy perspective. Our study establishes systematic methods of analysing embeddings in specialised domains. The current work considers QA task only. Future work may involve other tasks like summarization, or multi-modal settings.